OpenStudy (anonymous):

y=sinx-(1/2)sin2x for -2pi to 2pi find critical points and local max and minima

OpenStudy (mathmate):

Both are continuous functions, so the sum is continuous. Let f(x)=sinx-(1/2)sin2x f'(x)=cos(x)-cos(2x) => f'(x) is continuous and all points of f'(x) defined on the given interval. Solve for all x on the interval where f'(x)=0 to find the local extrema (and critical points). There are 7 solutions for f'(x)=0, including the two end-points, hence the same number of max-min.

OpenStudy (anonymous):

how do i find the asymptotes for this one

OpenStudy (mathmate):

There is none, both f(x) and f'(x) are continuous on the interval (in fact, from -inf to +inf).

OpenStudy (anonymous):

i only see 2 local max and 2 min in the interval am i missing somting?

OpenStudy (mathmate):

Are they symmetrical about x=0?

OpenStudy (anonymous):

maxima @ 2pi/3 and -4pi/3 right?

OpenStudy (anonymous):

minima @ 4pi/3 and -2pi/3

OpenStudy (anonymous):

do you count -2pi, 0 , and 2pi?

OpenStudy (mathmate):

How many points on the graph do you find f'(x)=0?

OpenStudy (mathmate):

Yes, those count as well, so there's your answer!

OpenStudy (mathmate):

However, check f"(x). If f"(x)=0, then it is NOT a max-min, just a critical point. There is more than one point of this kind.